d the marginal cost when erenti ate: (5 -2(3P- ) P) 2 t) =t. + 3t (t1)(t7) + 3 = 3x -
1 نقطة ()نقاط 3 السؤال The price paid for the rental of borrowed funds (usually expressed as a percentage of the rental of $100 per year) is commonly referred to as the inflation rate O aggregate price level O exchange rate interest rate O lai) alaäi 1 4 11a
d the marginal cost when erenti ate: (5 -2(3P- ) P) 2 t) =t. + 3t (t1)(t7) + 3 = 3x -
Answer #1

the tental ond dCommnyfeme aterest afe to as he fep(5F2) (a -) 3 3 to P (v)uvtov ;here U= 5P -2 V= 3P-1 11 dp15 3 IS 45P 5 b 3 fp 45P5 12F (45P- 1N-5 P 3 3t 3 f(t) ote crith 91e8pet to ftate)- () (CER+3t v-uv; bee U V2 ( V= Ceー)(セ)) +3 (at-) ((() (っ(ピッ) 2 ft3)t - (C49) (at t14t +3t ((th (##)) -32tt-a 4t+3+-l 2 fce3t-12+t*46-2it- 14t - (t-(t3+)2 2 4fx) 3x 2 P , * dx 30 C-a a-2) 2 3 3 2 3 2 2 3- (X-2) (x-2)2

Please rate if helpful. Comment if you have any doubt.